The raw data has been organised in the frequency table below.


games frequency


6 2


7 5


8 n


9 4


10 4


11 2


12 2


13 2


(a) Write down the value of n.


(b) Calculate the mean number of games played per set.


(c) What percentage of the sets had more than 10 games?


(d) What is the modal number of games?

1
Expert's answer
2022-01-25T16:50:35-0500

The numbers of games played in each set of a tennis tournament were



"9,7,8,11,9,6,10,8,12,6,8,13,7,""9,10,9,10,11,12,8,7,13,10,7,7"

The raw data has been organised in the frequency table below.



"\def\arraystretch{1.5}
 \begin{array}{c:c}
 games & frequency \\ \hline
 6 & 2\\
 \hdashline
 7 & 5\\
 \hdashline
 8 & n\\
 \hdashline
 9 & 4\\
 \hdashline
 10 & 4\\
 \hdashline
 11 & 2\\
 \hdashline
 12 & 2\\
 \hdashline
 13 & 2\\
 \hdashline
\end{array}"

(a)Β "n=4"


(b)



"mean=\dfrac{1}{2+5+4+4+4+2+2+2}\big(6(2)+7(5)""+8(4)+9(4)+10(4)+11(2)+12(2)+13(2)\big)=""=9.08"

(c)



"\dfrac{2+2+2}{25}\times100\%=24\%"

(d) The modal number of games is 7 (five times).